# Metadata-enhanced contrastive learning from retinal optical coherence tomography images

### v2

@@ 1 Introduction
Deep learning promises to revolutionise medical imaging by automating screening, diagnosis and monitoring of diseases (Davenport and Kalakota, 2019). Traditionally, the successful application of supervised deep learning has required large amounts of labelled data for training. This does not translate well to the medical domain where labelled images are scarce as their annotation typically requires the time and cost of a trained clinician. ImageNet, commonly used to train and benchmark algorithms for natural image processing, contains millions of labelled images (Deng et al., 2009), while the median size of datasets featured in the Medical Image Computing & Computer Assisted Intervention (MICCAI) conference in 2019 ranged from 120 to 180 subjects (Kiryati and Landau, 2021). There is a large disparity between the growing amount of medical data and the resources needed to annotate it for the research and development of deep learning tools in medicine (Willemink et al., 2020).
@@
Pretraining with contrastive frameworks enables deep learning models to reach the same level of performance using fewer labelled training samples (Chen et al., 2020b). They work by contrasting two augmented views of each image per batch, assuming their similarity depends on whether or not they originate from the same image. However, they are prone to provide misleading learning signals when the true set of inter-image relationships are unknown. As illustrated in (Khosla et al., 2020), removing negative pairs of images from the same ImageNet class improves downstream performance. In contrast, large volumes of medical images are not typically annotated with gold-standard class labels for disease stage or type, but are commonly accompanied by metadata such as the date of the medical scan and the patient’s anonymised identity. This widely available information has the potential to indicate the true set of inter-image relationships used in contrastive frameworks in the medical domain.
@@
Optical coherence tomography (OCT) enables low-cost, non-invasive imaging of the eye. This has led to the accumulation of large, metadata-enriched retinal datasets that approach the size of those for natural images. Hence, ophthalmology is ideally suited for adapting advances in self-supervised learning to the medical domain. A particularly challenging task is the diagnosis and prognosis of age-related macular degeneration (AMD) in retinal OCT. AMD, the leading cause of irreversible blindness in the elderly, is a progressive disease and is that is projected to increase in prevalence by nearly 50% from 196 million cases worldwide in 2020 to 288 million by 2040 (Wong et al., 2014).
@@
In this paper we extend contrastive pretraining with strategies leveraging time series of widely available metadata in longitudinal retinal imaging datasets (see Figure 1). We conduct our analysis across seven diverse downstream tasks related to AMD in two large datasets totalling 170,427 OCT images of 7,912 patients collected in the scope of the PINNACLE study (Sutton et al., 2022). Overall, our contributions and key findings include:
@@ 2.1 Self-supervised learning
Self-supervised pretraining has gained momentum in recent years as a method for creating generalisable representations of unlabelled data. Pretrained models can then be finetuned on supervised tasks resulting in faster training and requiring fewer annotations to reach good performance. It works by solving so-called pretext tasks that create learning signals without human supervision. Initially pretext tasks were derived from ad-hoc heuristics such as predicting image rotation (Komodakis and Gidaris, 2018) and colourisation (Zhang et al., 2016) but these were surpassed by more general context matching tasks such as solving jigsaw puzzles (Noroozi and Favaro, 2016) and predicting one local context from another (Oord et al., 2018).
Contrastive frameworks (Chen et al., 2020a; He et al., 2020; Grill et al., 2020; Chen and He, 2021) currently advance the state-of-the-art in self-supervised learning. In their seminal work, Chen et al. 2020a proposed SimCLR which creates a context matching task by combining a set of image transformations with a contrastive loss. They define similarity at the image level, such that two augmented views of the same image create a positive contrastive pair and views from different images make negative pairs. Models then learn to maximise similarity in representation space between positive pairs, and minimise similarity between negative pairs. They find that a diverse set of transformations, including random crops, rotations and colour shifts, helps in building more generalisable representations than could be learned from any one pretext task.
Khosla et al. 2020 show that standard contrastive learning on ImageNet is degraded by misleading negative pairs consisting of images belonging to the same class, and restore performance by instead making these pairs positive. In Bootstrap Your Own Latent (BYOL) Grill et al. 2020 remove negative pairs altogether and employ imbalanced student and teacher networks to prevent a degenerate solution. Overall, choosing an appropriate and diverse set of transformations as well as the correct handling of false negative pairs are crucial for the successful application of contrastive learning.
@@ 2.2 Self-supervised learning in the medical domain
Early applications of self-supervised learning to medical images also relied on ad-hoc pretext tasks such as image registration (Li and Fan, 2017) or region-specific tasks such as locating anatomical landmarks in the heart (Bai et al., 2019), distance between random patches in the brain (Spitzer et al., 2018) and estimating retinal thickness in the eye (Holmberg et al., 2020). Chen et al. 2019 use a generic context restoration task to improve plane detection in ultrasound, landmark detection in CT and tumour segmentation in MRI.
There are very few applications of standard contrastive learning to medical datasets (Ghesu et al., 2022). Many transformations used to generate views of natural images are not compatible with medical images. In contrast to colour images medical scans are often single-channel rendering two of the most performant augmentations in SimCLR, colour shift and greyscale, inapplicable. Additionally, medical scans often have a strong prior on their content. For example, studies typically collect scans targeting a single anatomical region or medical condition. Views generated from different images that feature similar disease-related or anatomical content will nonetheless form misleading negative pairs and degrade the learning signal.
To address these problems others have used domain specific information. Chen et al. 2021 argue that views generated from any two frames of the same ultrasound video are highly correlated. By switching their contrastive relationship from negative to positive they show improvement in detection of COVID-19 and pneumonia. Others generate more informative positive pairs by choosing bilateral views to improve classification of conditions in dermatological photographs (Azizi et al., 2021), multi-modal views between genetic information and fundus images for classifying cardiac disease (Taleb et al., 2022), views from the same temporal longitude for improving lung segmentation (Zeng et al., 2021a) and views with the same disease label for classifying pleural effusion in chest X-ray (Vu et al., 2021). Two studies create positive pairs from spatially proximal 2D slices to improve volumetric segmentation of cardiac MRI (Zeng et al., 2021b; Chaitanya et al., 2020). Ciga et al. 2022 limit negative pairs to histopathological images with different staining and tissue types for improved detection of cancer.
@@ 2.3 Deep learning for retinal imaging and AMD
Most existing applications of deep learning to retinal images are fully supervised (Bogunović et al., 2017; Lee et al., 2017; Schmidt-Erfurth et al., 2018; De Fauw et al., 2018). There have been very few applications of self-supervised learning to retinal images. Holmberg et al. 2020 use a multi-modal approach to improve detection of retinopathy by pretraining models to use fundus photographs to predict retinal thickness maps derived from OCT. Srinivasan et al. 2021 find that contrastive pretraining on ImageNet boosted downstream analysis of diabetic retinopathy in small fundus imaging datasets.
Self-supervised pretraining for AMD has so far included one work using fundus photographs (Yellapragada et al., 2022) and another employing a temporal ordering task to learn late AMD features from longitudes in OCT (Rivail et al., 2019). More recently (Zhou et al., 2023) released RETFound, a foundation model for retinal images trained on over 700,000 OCT images. However, there are no studies comprehensively testing the potential role of metadata for improving self-supervised learning in the domain of retinal OCT.
@@ 3.1 OCT datasets
In this paper we use two datasets collected as part of the PINNACLE project (Sutton et al., 2022). They each contain longitudinal data consisting of retinal OCT scans of patients with AMD. For both datasets images were acquired using Topcon 3D OCT devices (Topcon Corporation, Tokyo, Japan). The first dataset, collected at the Southampton Eye Unit, consists of 48,825 OCT scans and followed 6,368 eyes from 3,498 patients. A subset of 2,037 images are healthy control scans curated from 570 patients. In Figure 3a we show the number of times eyes were scanned. The average eye was scanned 7.7 times over an average duration of 1.9 years, with longitudes featuring AMD spanning 2.3 years. The second dataset, collected at Moorfields Eye Hospital, is larger containing 121,602 OCT scans of 7,336 eyes from 3,844 patients. The longitudinal dimension is also greater, with longitudes containing on average 16.6 scans spanning 3.5 years. We extracted the mediolateral 2D slice centred at the fovea and resampled to 208$\times$256 pixels with a pixel size of 7.0×23.4 $\mu m^{2}$, half the median resolution. We elect to maintain these as separate datasets to test whether any benefits from our pretraining methods generalise to different hospitals and collection sites.
@@
AMD stage labels, indicating whether a patient has early or late stage AMD, were extracted automatically from the electronic health records. For 39% of scans it was possible to extract an AMD stage label in this manner, and the remaining 61% were left unlabelled. Late stage AMD was further characterised into either wet (choroidal neovascularization) or dry (geographic atrophy) AMD. Clinical information also included visual acuity, a metric for assessing functional quality of vision, which measured the logarithm of the minimum angle of resolution (LogMAR) perceptible to the patient. We rescale this measurement via a linear mapping to find the Letter Score, indicating how many letters, from 0 to 95, can be read on a letter chart (Beck et al., 2003). Finally, we obtained segmentations of eleven retinal layers using the Iowa Reference Algorithms (Retinal Image Analysis Lab, Iowa Institute for Biomedical Imaging, USA), a graph-cut-based segmentation algorithm (Li et al., 2005; Garvin et al., 2009; Abràmoff et al., 2010).
@@ 3.2 Metadata-enhanced contrastive learning
In this study, we first extend SimCLR (Chen et al., 2020a) as it is the most widely adopted and paradigmatic contrastive method (Chaitanya et al., 2020; Azizi et al., 2021; Ciga et al., 2022; Taleb et al., 2022). For a second standard baseline we include BYOL (Grill et al., 2020) for its exclusive use of positive pairs. As alluded to in section 2.2 standard contrastive learning suffers from two issues in medical applications. The first is that several transformations used by SimCLR and BYOL, such as hue, saturation and colour dropping, are not applicable in single-channel medical scans. The second is that many negative contrastive pairs will in fact feature highly similar views since all OCT images depict the same anatomy and limited set of pathologies, especially if both images originate from the same eye or patient.
@@ 3.3.1 Contrastive pretraining protocol
Both standard and metadata-enhanced contrastive pretraining variants, for both SimCLR, BYOL, use a ResNet50 (4x) backbone with half precision initialised according to the He protocol (He et al., 2016). In both datasets we allocate 4,800 samples for validation, 4,800 for testing and the remainder for pretraining. Models then train with the AdamW optimiser using a learning rate of $5\cdot 10^{-4}$ and momentum of $0.9$ for 120,000 steps. We use linear warmup for the first 1,200 steps and decay the learning rate with the cosine schedule without restarts. The batch size is 384 and we use the model after the final training step for downstream evaluation. Additional details are listed in A.2. Due to the sensitivity of contrastive frameworks to the chosen transformations we closely follow those used by BYOL (Grill et al., 2020), transferring as many as are applicable to retinal OCT. We first, with probability 0.8, augment each scan by varying image brightness and contrast using maximum relative changes of 0.4. Then we rotate by a random angle up to 8 degrees, crop centrally to resolution 188$\times$236 to remove rotation border artefacts, flip horizontally with probability 0.5, randomly crop to a scale between 0.25 and 1.0 with aspect ratio between 3/4 and 4/3, and finish by resizing to 192$\times$192. We chose not to use Gaussian blur as we, in accordance with (Azizi et al., 2021), found it obfuscated local textural features in the retinal images. We also omit hue, saturation and colour dropping as these augmentations do not have any effect on greyscale images.
@@ 3.3.2 Pretrained baselines
To fairly assess the benefit of metadata-enhanced learning we compare against three pretrained baselines. The first baseline is a model pretrained by (Chen et al., 2020a) in SimCLR on natural images in ImageNet. Our second baseline uses a more medically aligned model pretrained to classify 165 different conditions on RadImageNet (Mei et al., 2022). We also compare our methods against RETFound, a foundation model for retinal images trained on over 700,000 OCT images (Zhou et al., 2023). RETFound was trained as a Masked Autoencoder (He et al., 2022) and uses a ViT-Large backbone. We use the version specifically trained on OCT data, which was found to be highly performant in retinal disease prediction, including AMD, in external datasets. These baselines, pretrained on natural images, medical images and external retinal OCT images, offer increasing levels of in-domain pretraining against which to compare in-dataset contrastive learning, and our metadata-enhanced extensions.
@@ 3.3.3 Finetuning protocol
To compare different pretrained models we adopt the standard linear evaluation finetuning protocol used by (Chen et al., 2020a). We first freeze the pretrained encoder and replace any dense projection layers used in pretraining with a single linear layer that projects from the latent space to the dimension of the supervised label. More specifically, for the segmentation task this involves a 11x192 dimension matrix and for all other problems the output is size 1. For classification problems we then apply a sigmoid activation function and for segmentation and regression problems we use $\tanh$ activation and scale labels to the range [-1,1]. We use a batch size of 1024 and ensure we use the same train, validation and test splits as for pretraining. We then train all models for 3000 epochs using the early stopping if downstream validation performance does not improve for 600 steps. To fit the model we use an AdamW optimiser with momentum of $0.9$, weight decay of $1\cdot 10^{-2}$ and a learning rate of $5\cdot 10^{-4}$ using the cosine annealing schedule after 50 linear warmup epochs. These hyperparameters are the same as were used to evaluate RETFound in (He et al., 2022). Moreover, for all models pretrained on external dataset we also resize the input image to the size they were trained on, which was 224x224. In training we employ a weaker form of the augmentations described in the pretraining protocol. Finally, we report performance on the test set of the model checkpoint with the best validation downstream performance.

### v3

@@ 1 Introduction
Deep learning promises to revolutionise medical imaging by automating screening, diagnosis and monitoring of diseases [14]. Traditionally, the successful application of supervised deep learning has required large amounts of labelled data for training. This does not translate well to the medical domain where labelled images are scarce as their annotation typically requires the time and cost of a trained clinician. ImageNet, commonly used to train and benchmark algorithms for natural image processing, contains millions of labelled images [16], while the median size of datasets featured in the Medical Image Computing & Computer Assisted Intervention (MICCAI) conference in 2019 ranged from 120 to 180 subjects [27]. There is a large disparity between the growing amount of medical data and the resources needed to annotate it for the research and development of deep learning tools in medicine [45].
@@
Pretraining with contrastive frameworks enables deep learning models to reach the same level of performance using fewer labelled training samples [9]. They work by contrasting two augmented views of each image per batch, assuming their similarity depends on whether or not they originate from the same image. However, they are prone to provide misleading learning signals when the true set of inter-image relationships are unknown. As illustrated in [26], removing negative pairs of images from the same ImageNet class improves downstream performance. In contrast, large volumes of medical images are not typically annotated with gold-standard class labels for disease stage or type, but are commonly accompanied by metadata such as the date of the medical scan and the patient’s anonymised identity. This widely available information has the potential to indicate the true set of inter-image relationships used in contrastive frameworks in the medical domain.
@@
Optical coherence tomography (OCT) enables low-cost, non-invasive imaging of the eye. This has led to the accumulation of large, metadata-enriched retinal datasets that approach the size of those for natural images. Hence, ophthalmology is ideally suited for adapting advances in self-supervised learning to the medical domain. A particularly challenging task is the diagnosis and prognosis of age-related macular degeneration (AMD) in retinal OCT. AMD, the leading cause of irreversible blindness in the elderly, is a progressive disease and is that is projected to increase in prevalence by nearly 50% from 196 million cases worldwide in 2020 to 288 million by 2040 [46].
@@
In this paper we extend contrastive pretraining with strategies leveraging time series of widely available metadata in longitudinal retinal imaging datasets (see Figure 1). We conduct our analysis across seven diverse downstream tasks related to AMD in two large datasets totalling 170,427 OCT images of 7,912 patients collected in the scope of the PINNACLE study [41]. Overall, our contributions and key findings include:
@@ 2.1 Self-supervised learning
Self-supervised pretraining has gained momentum in recent years as a method for creating generalisable representations of unlabelled data. Pretrained models can then be finetuned on supervised tasks resulting in faster training and requiring fewer annotations to reach good performance. It works by solving so-called pretext tasks that create learning signals without human supervision. Initially pretext tasks were derived from ad-hoc heuristics such as predicting image rotation [28] and colourisation [50] but these were surpassed by more general context matching tasks such as solving jigsaw puzzles [34] and predicting one local context from another [35].
Contrastive frameworks [8, 21, 19, 10] currently advance the state-of-the-art in self-supervised learning. In their seminal work, Chen et al. 8 proposed SimCLR which creates a context matching task by combining a set of image transformations with a contrastive loss. They define similarity at the image level, such that two augmented views of the same image create a positive contrastive pair and views from different images make negative pairs. Models then learn to maximise similarity in representation space between positive pairs, and minimise similarity between negative pairs. They find that a diverse set of transformations, including random crops, rotations and colour shifts, helps in building more generalisable representations than could be learned from any one pretext task.
Khosla et al. 26 show that standard contrastive learning on ImageNet is degraded by misleading negative pairs consisting of images belonging to the same class, and restore performance by instead making these pairs positive. In Bootstrap Your Own Latent (BYOL) Grill et al. 19 remove negative pairs altogether and employ imbalanced student and teacher networks to prevent a degenerate solution. Overall, choosing an appropriate and diverse set of transformations as well as the correct handling of false negative pairs are crucial for the successful application of contrastive learning.
@@ 2.2 Self-supervised learning in the medical domain
Early applications of self-supervised learning to medical images also relied on ad-hoc pretext tasks such as image registration [30] or region-specific tasks such as locating anatomical landmarks in the heart [3], distance between random patches in the brain [39] and estimating retinal thickness in the eye [24]. Chen et al. 7 use a generic context restoration task to improve plane detection in ultrasound, landmark detection in CT and tumour segmentation in MRI.
There are very few applications of standard contrastive learning to medical datasets [18]. Many transformations used to generate views of natural images are not compatible with medical images. In contrast to colour images medical scans are often single-channel rendering two of the most performant augmentations in SimCLR, colour shift and greyscale, inapplicable. Additionally, medical scans often have a strong prior on their content. For example, studies typically collect scans targeting a single anatomical region or medical condition. Views generated from different images that feature similar disease-related or anatomical content will nonetheless form misleading negative pairs and degrade the learning signal.
To address these problems others have used domain specific information. Chen et al. 11 argue that views generated from any two frames of the same ultrasound video are highly correlated. By switching their contrastive relationship from negative to positive they show improvement in detection of COVID-19 and pneumonia. Others generate more informative positive pairs by choosing bilateral views to improve classification of conditions in dermatological photographs [2], multi-modal views between genetic information and fundus images for classifying cardiac disease [42], views from the same temporal longitude for improving lung segmentation [48] and views with the same disease label for classifying pleural effusion in chest X-ray [43]. Two studies create positive pairs from spatially proximal 2D slices to improve volumetric segmentation of cardiac MRI [49, 6]. Ciga et al. 13 limit negative pairs to histopathological images with different staining and tissue types for improved detection of cancer.
@@ 2.3 Deep learning for retinal imaging and AMD
Most existing applications of deep learning to retinal images are fully supervised [5, 29, 37, 15]. There have been very few applications of self-supervised learning to retinal images. Holmberg et al. 24 use a multi-modal approach to improve detection of retinopathy by pretraining models to use fundus photographs to predict retinal thickness maps derived from OCT. Srinivasan et al. 40 find that contrastive pretraining on ImageNet boosted downstream analysis of diabetic retinopathy in small fundus imaging datasets.
Self-supervised pretraining for AMD has so far included one work using fundus photographs [47] and another employing a temporal ordering task to learn late AMD features from longitudes in OCT [36]. More recently [51] released RETFound, a foundation model for retinal images trained on over 700,000 OCT images. However, there are no studies comprehensively testing the potential role of metadata for improving self-supervised learning in the domain of retinal OCT.
@@ 3.1 OCT datasets
In this paper we use two datasets collected as part of the PINNACLE project [41]. They each contain longitudinal data consisting of retinal OCT scans of patients with AMD. For both datasets images were acquired using Topcon 3D OCT devices (Topcon Corporation, Tokyo, Japan). The first dataset, collected at the Southampton Eye Unit, consists of 48,825 OCT scans and followed 6,368 eyes from 3,498 patients. A subset of 2,037 images are healthy control scans curated from 570 patients. In Figure 3a we show the number of times eyes were scanned. The average eye was scanned 7.7 times over an average duration of 1.9 years, with longitudes featuring AMD spanning 2.3 years. The second dataset, collected at Moorfields Eye Hospital, is larger containing 121,602 OCT scans of 7,336 eyes from 3,844 patients. The longitudinal dimension is also greater, with longitudes containing on average 16.6 scans spanning 3.5 years. We extracted the mediolateral 2D slice centred at the fovea and resampled to 208$\times$256 pixels with a pixel size of 7.0×23.4 $\mu m^{2}$, half the median resolution. We elect to maintain these as separate datasets to test whether any benefits from our pretraining methods generalise to different hospitals and collection sites.
@@
AMD stage labels, indicating whether a patient has early or late stage AMD, were extracted automatically from the electronic health records. For 39% of scans it was possible to extract an AMD stage label in this manner, and the remaining 61% were left unlabelled. Late stage AMD was further characterised into either wet (choroidal neovascularization) or dry (geographic atrophy) AMD. Clinical information also included visual acuity, a metric for assessing functional quality of vision, which measured the logarithm of the minimum angle of resolution (LogMAR) perceptible to the patient. We rescale this measurement via a linear mapping to find the Letter Score. This is a more intuitive metric which indicates how many letters, from 0 to 95, could be read by the patient on a letter chart [4]. Finally, we obtained segmentations of eleven retinal layers using the Iowa Reference Algorithms (Retinal Image Analysis Lab, Iowa Institute for Biomedical Imaging, USA), a graph-cut-based segmentation algorithm [32, 17, 1].
@@ 3.2 Metadata-enhanced contrastive learning
In this study, we first extend SimCLR [8] as it is the most widely adopted and paradigmatic contrastive method [6, 2, 13, 42]. For a second standard baseline we include BYOL [19] for its exclusive use of positive pairs. As alluded to in section 2.2 standard contrastive learning suffers from two issues in medical applications. The first is that several transformations used by SimCLR and BYOL, such as hue, saturation and colour dropping, are not applicable in single-channel medical scans. The second is that many negative contrastive pairs will in fact feature highly similar views since all OCT images depict the same anatomy and limited set of pathologies, especially if both images originate from the same eye or patient.
@@ 3.3.1 Contrastive pretraining protocol
Both standard and metadata-enhanced contrastive pretraining variants, for both SimCLR, BYOL, use a ResNet50 (4x) backbone with half precision, and were randomly initialised before training according to the He protocol [22]. In both datasets we allocate 4,800 samples for validation, 4,800 for testing and the remainder for pretraining. Models then train with the AdamW optimiser using a learning rate of $5\cdot 10^{-4}$ and momentum of $0.9$ for a fixed number of 120,000 optimisation steps. We use linear warmup for the first 1,200 steps and decay the learning rate with the cosine schedule without restarts. The batch size is 384 and we use the model after the final training step for downstream evaluation. Additional details are listed in A.2. Due to the sensitivity of contrastive frameworks to the chosen transformations we closely follow those used by BYOL [19], transferring as many as are applicable to retinal OCT. We first, with probability 0.8, augment each scan by varying image brightness and contrast using maximum relative changes of 0.4. Then we rotate by a random angle up to 8 degrees, crop centrally to resolution 188$\times$236 to remove rotation border artefacts, flip horizontally with probability 0.5, randomly crop to a scale between 0.25 and 1.0 with aspect ratio between 3/4 and 4/3, and finish by resizing to 192$\times$192. We chose not to use Gaussian blur as we, in accordance with [2], found it obfuscated local textural features in the retinal images. We also omit hue, saturation and colour dropping as these augmentations do not have any effect on greyscale images.
@@ 3.3.2 Pretrained baselines
To fairly assess the benefit of metadata-enhanced learning we compare against three pretrained baselines. The first baseline is a model pretrained by [8] in SimCLR on natural images in ImageNet. Our second baseline uses a more medically aligned model pretrained to classify 165 different conditions on RadImageNet [33]. We also compare our methods against RETFound, a foundation model for retinal images trained on over 700,000 OCT images [51]. RETFound was trained as a Masked Autoencoder [20] and uses a ViT-Large backbone. We use the version specifically trained on OCT data, which was found to be highly performant in retinal disease prediction, including AMD, in external datasets. These baselines, pretrained on natural images, medical images and external retinal OCT images, offer increasing levels of in-domain pretraining against which to compare in-dataset contrastive learning, and our metadata-enhanced extensions.
@@ 3.3.3 Finetuning protocol
To compare different pretrained models we adopt the standard linear evaluation finetuning protocol used by [8]. We first freeze the pretrained encoder and replace any dense projection layers used in pretraining with a single linear layer that projects from the latent space to the dimension of the supervised label. More specifically, for the segmentation task this involves a 11x192 dimension matrix and for all other problems the output is size 1. For classification problems we then apply a sigmoid activation function and for segmentation and regression problems we use $\tanh$ activation and scale labels to the range [-1,1]. We use a batch size of 1024 and ensure we use the same train, validation and test splits as for pretraining. We then train all models for 3000 epochs using the early stopping if downstream validation performance does not improve for 600 steps. To fit the model we use an AdamW optimiser with momentum of $0.9$, weight decay of $1\cdot 10^{-2}$ and a learning rate of $5\cdot 10^{-4}$ using the cosine annealing schedule after 50 linear warmup epochs. These hyperparameters are the same as were used to evaluate RETFound in [20]. Moreover, for all models pretrained on external dataset we also resize the input image to the size they were trained on, which was 224x224. In training we employ a weaker form of the augmentations described in the pretraining protocol. Finally, we report performance on the test set of the model checkpoint with the best validation downstream performance.